OpenStudy (anonymous):

What mass of Mg is formed by passing a current of 5.00A through molten MgCl2 for 4.90 days? Can someone help me with this please?

OpenStudy (jfraser):

you need to use the relationships of \[current = \frac{charge}{time}\] and \[\frac{charge}{mole \space of \space electrons}\] to start with time (4.9 days, which must be converted into seconds) into mass of Mg

OpenStudy (jfraser):

as an example: if a current of 2.5 amps is passed through molten AlCl3 for one day, the mass of aluminum would be:\[8.64*10^5 \space seconds * (\frac{2.5 C}{1 \sec})*(\frac{1mole \space electrons}{96,500C})*(\frac{1mol \space Al}{3mol \space electrons})*(\frac{27g \space Al}{1mol \space Al})\] \[=20.1g \space Al\]

OpenStudy (jfraser):

the starting time must be in seconds, because current in amps is coulombs per second 96,500C per mole of electrons is faraday's constant, learn it 3moles of electrons will fully reduce 1 mole of Al from \(Al^{+3}\) to \(Al^0\) and 27g/mol is the molar mass of Al

OpenStudy (anonymous):

Faraday is actually 96, 485 C/mole electrons